OpenStudy (strawberry17):

How would you solve this: Mark is graduating from college in six months, but he will need a loan in the amount of $7,038 for his last semester. He may either receive an unsubsidized Stafford Loan with an interest rate of 6.8%, compounded monthly, or his parents may obtain a PLUS Loan with an interest rate of 7.8%, compounded monthly. The Stafford Loan has a grace period of six months from the time of graduation. Which loan will have a higher balance at the time of repayment and by how much?

OpenStudy (strawberry17):

The PLUS Loan has a higher balance by $214.80. The PLUS Loan has a higher balance by $690.26. The Stafford Loan has a higher balance by $184.50. The Stafford Loan has a higher balance by $259.74.

OpenStudy (strawberry17):

This is how I solve it: Stafford Loan B = P(1+i)^nt B = 7038(1 + .068/12)^12• 12/12 B = 7531.79

OpenStudy (strawberry17):

PLUS Loan B =P(1+i)^nt B = 7038(1+ .078/12)12• 6/12 B = 7,316.98

OpenStudy (strawberry17):

7,531.79 - 7,316.98 = 214.81 This is saying that the Stafford Loan is higher by 214.81, but the answer has it the opposite.

OpenStudy (strawberry17):

Did I solve it right?

OpenStudy (strawberry17):

@jim_thompson5910 can you help me please?

jimthompson5910 (jim_thompson5910):

one sec, still thinking

jimthompson5910 (jim_thompson5910):

You did everything right and the stafford loan has a higher balance because an extra 6 months went by...just not sure how they think it's the PLUS loan that has the higher balance

jimthompson5910 (jim_thompson5910):

I want to say it's a typo, but I might be missing something

OpenStudy (strawberry17):

thanks

OpenStudy (strawberry17):

I just needed someone who knew to check

jimthompson5910 (jim_thompson5910):

personally I would go with choice A and hope they just made a silly error

OpenStudy (strawberry17):

k